[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue Aug 16, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Tue Aug 16 11:33:34 UTC 2016


HSU Shift Log for Tue Aug 16 2016 07:33:28 GMT-0400 (EDT)

Shift: H

Forecaster: L Avila

Technical Issues Summary:

ATCF issues: A 13 is not in the ATCF as in the previous shift.

N-AWIPS/AWIPS2 issues:

Other Techincal issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160816/74f2cd8b/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list